Find the equation of the line through point (1,2) and parallel to 3x+4y=12. Use a slash ("/") for fractions (e.g. 1/2 for 12). DO not include spaces in your answers.
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

Find the equation of the line through point (1,2) and parallel to 3x+4y=12.
------------
Find the slope, m, f the given line.
---
Use y-y1 = m*(x-x1) for the equation
(x1,y1) is the point (1,2)

Answer by greenestamps(13200)



Any line parallel to 3x+4y=12 has an equation of the form 3x+4y=C for some constant C.

Since the point (1,2) is on the line,

3x+4y = 3(1)+4(2) = 3+8 = 11

So the equation is 3x+4y=11.